Two jets leave an air base at the same time and travel in opposite directions. One jet travels 15mph slower than the other. If the two jets are 6730 miles apart after 5 hours, what is the rate of each jet?
Answer by ikleyn(52814)

.

Let r be the rate of the slower jet;  then the rate of the faster jet is (r+15) mph.


The total separating distance is

    5r + 5*(r+15),


and it is equal to 6730.  So your base equation is

    5r + 5*(r+15) = 6730.


Simplify and find r

    5r + 5r + 75 = 6730

    10r          = 6730 - 75 = 6655

      r                      = 665.5.


ANSWER.  The rate of the slower jet is  665.5 mph.

         The rate of the faster jet is  665.5 + 15 = 680.5 mph.


CHECK.  5*665.5 + 5*680.5 = 6730  separating distance in 5 hours.   ! Correct !

Solved.

-----------------

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.